# Spin excitations of the Shastry-Sutherland model – altermagnetism and proximate deconfined quantum criticality

### v1

@@ Abstract
Symmetry plays a crucial role in condensed matter physics. In quantum magnetism, it dictates a number of exotic phenomena, including deconfined quantum criticality and altermagnetism. Here, by studying the spin excitations of the $S=1/2$ antiferromagnetic Shastry-Sutherland model, we show that the Néel antiferromagnetic state in this model is an altermagnet featuring a non-relativistic splitting between two chiral magnon bands. Moreover, we identify a Higgs mode in the longitudinal excitation channel, whose gap softens when approaching the antiferromagnetic to plaquette valence bond solid transition, implying the appearance of nearly deconfined excitations. However, the splitting between the two magnon bands (Goldstone modes) remains finite at the transition. These results indicate that the transition is weakly first-order and proximate to a putative deconfined quantum critical point. We find that the altermagnetism provides a sensitive means to probe the deconfined quantum criticality.
@@
Introduction.—Symmetry is essential in physics. It provides an organizational principle to classify complex states of matter, and based on this, many exotic phenomena are understood TI_Review ; Senthil_AnnRevCondMattPhys_2015 ; Witten_NP_2018 . One prominent example is the recently proposed altermagnet, a collinear antiferromagnetic (AFM) state in which the two sublattices with oppositely polarized spins are connected by the rotational symmetry instead of inversion or translation Smejkal_PRX_2022A ; Smejkal_PRX_2022B . The distinct symmetry, which is organized by the spin space group theory Yang_arXiv:2105.12738 , causes a non-relativistic spin-dependent splitting of energy bands and may give rise to some exotic effects, such as anomalous Hall effect Smejkal_PRX_2022A ; Smejkal_NatRevMater7_2022 . Up to now, most studies focus on itinerant systems Gonzalez_Betancourt_PRL_2023 ; Hariki_PRL_2024 ; Ferrari_arXiv:2408.00841 . For insulating altermagnets, though a symmetry dictated splitting in magnon bands has been predicted Smejkal_PRL131_256703_2023 , existing studies are mostly limited to several model systems Ma_PRB110_064426_2024 ; Weissenhofer_PRB110_094427_2024 ; Yao_arXiv_2024 ; Zhaojize_arxiv .
@@
Symmetry can also dictate phase transitions. In the Landau-Ginzburg-Wilson paradigm, a system transitions from a disordered state to an ordered one by spontaneously breaking a symmetry. With quantum fluctuations the transition can be continuous and occur at zero temperature, giving rise to intriguing quantum critical phenomena sachdev_2011 ; Coldea_2010 ; Senthil2004 ; Cui2023DQCP ; Cui2019SCVO ; Lee2019 ; E8_2021 ; Xu2022 ; agrawal2020quantum ; fey2019quantum ; zhang2011exploring . For a transition between two ordered phases, the Landau paradigm predicts a first-order transition. However, this scenario has recently been challenged. It was proposed that the transition between a valence bond solid (VBS) and an antiferromagnetic (AFM) phase can be continuous, e.g., via a deconfined quantum critical point (DQCP) Senthil2004 . At the DQCP, the establishment of one order is accompanied by the destruction of the other order via proliferation of topological excitations, giving rise to emergent deconfined fractionalized excitations as well as enhanced continuous symmetry among the seemingly unrelated two types of order parameters. This scenario has motivated extensive studies on DQCP Sandvik2007 ; Meng_EasyplaneDQCP ; Shao_Science ; Senthil_Vojta_Sachdev:2004 ; Liu_Vojta:2022 ; Guo_PRL_2024 . However, whether a DQCP can be realized in quantum spin models is still an open question.
@@
Tuning the frustration is a promising way to induce a VBS-to-AFM transition, and a prominent example is given by the Shastry-Sutherland (SS) model Shastry1981 . The model is defined on the SS lattice as illustrated in Fig. 1(a). It contains orthogonal spin dimers (with coupling $J^{\prime}$) connected by frustrated nearest neighbor (n.n.) bonds (with coupling $J$). With increasing $J/J^{\prime}$, the ground state experiences a series of quantum phase transitions from a dimer singlet phase to a plaquette VBS at $J/J^{\prime}\approx 0.68$, then to a Néel AFM state for $J/J^{\prime}\sim 0.8$ (see Fig. 1(b)).Koga2000 ; Chung2001 ; Pixley2014 ; Corboz2013 ; Boos2019 ; Lee2019 ; yang2021quantum ; Sengupta_PRL:2013 ; Wang_PRL_2023 . Although the dimer-to-plaquette transition is clearly strongly first-order, the nature of the plaquette-to-AFM transition remains elusive: Several tensor network calculations Corboz2013 ; Xi2023 showed it to be weakly first-order, while some other study yang2021quantum found an intervening gapless quantum spin liquid in between the plaquette and AFM phases. Interestingly, one DMRG work Lee2019 suggested the transition is via a DQCP with an emergent $O(4)$ symmetry, and quantum scaling behavior near this point has been revealed by a finite-size tensor network analysis WenyuanLiu_PRL_2024 . One prominent signature of a DQCP is the emergence of deconfined fractionalized spin excitations. Although spin excitations in both plaquette and AFM states have been extensively studied Mila_BOT ; Pinaki_PRB92_094440_2015 ; WangZ2018 ; Wang2023 , whether deconfined excitations emerge at the transition is still unclear.
@@
The phase diagram of the SS model well describes the evolution of low-temperature phases of the quasi-2D antiferromagnet $\mathrm{SrCu}_{2}(\mathrm{BO}_{3})_{2}$ under pressure tuning Kageyama1999 ; PhysRevLett.82.3701 ; Koga2000 ; Waki2007 ; haravifard2016crystallization ; Zayed2017 ; Bettler2020 ; Guo2020 ; Jimenez2021 , giving the hope to realize a DQCP in this material. Remarkably, a recent NMR study in this compound Cui2023DQCP reported an unusually large anomalous dimension $\eta$ at a field-induced plaquette-to-AFM transition under pressure, implying behavior in proximity to a DQCP. However, direct spectral evidence on the deconfined excitations at zero field is absent.
In this Letter, we study spin excitations of the SS model by using the state-of-the-art tensor network method for infinite lattice systems. This prevents from the severe finite-size effect in usual numerical calculations and offers a new perspective on the nature of emergent DQCP phenomena. We show that the altermagnetic nature of the AFM state causes a finite splitting with opposite chirality in the magnon bands along the $\Gamma$ to M direction of the Brillouin zone. Besides the transverse excitations, we also identify a Higgs mode in the longitudinal excitation channel, whose energy decreases with reducing $J/J^{\prime}$ ratio. Approaching the plaquette-to-AFM transition, the gaps of the Higgs mode and the triplet excitations in the plaquette phase both rapidly drop close to zero, implying the emergence of nearly deconfined excitations. However, the splitting of the magnon bands remains finite at the transition. These results suggest the transition is weakly first-order, in proximity to a DQCP with an emergent O(4) symmetry. In general, we show that the split magnon bands caused by the altermagnetism serve as a sensitive probe for a deconfined quantum phase transition.
@@
Figure 1: (a): Sketch of the SS lattice and the corresponding altermagnetism. The two sublattices of the Néel state are connected by a $C_{4}$ rotation of the lattice about the center of the empty plaquette (without the dimer bond). A time-reversal symmetry then recovers the Néel order. (b): Sketched phase diagram of the SS model with $J/J^{\prime}$. The $J/J^{\prime}$ values at the two transitions are taken from Ref. Xi2023 . (c): LSW dispersion in the AFM state of the SS model showing the chiral magnon bands with a non-zero splitting along the M-$\Gamma$-M direction of the Brillouin zone. $\mathcal{L}$($\mathcal{R}$) refers to the left(right)-handed chiral magnon bands.
Model and method.—The Hamiltonian of the SS model reads as
@@
[…]
Figure 2: (a): Spin DSF in the AFM phase of the SS model calculated by the LSW theory. The inset shows high symmetric directions of the Brillouin zone. (b): Total spin DSF ($\mathcal{S}^{xx+yy+zz}$) in the AFM phase calculated by the iPEPS method with $D=3$ at $J/J^{\prime}=1.1$. (c): The transverse component ($\mathcal{S}^{xx+yy}$) of the results in (b). (d): The longitudinal component ($\mathcal{S}^{zz}$) of the results in (b). The dashed lines in (b) and (d) label the Higgs mode.
@@
Altermagnetism and chiral magnons in the SS model.—The space group of the SS lattice is non-symmophic and the two sublattices of the Néel AFM state are connected by neither translational nor inversion, but a $C_{4}$ rotation about the center of the empty plaquette. The AFM order then recovers by further applying a time-reversal $\mathcal{T}$ symmetry to the spins (see Fig. 1(a)). This indicates that the Néel AFM state on the SS lattice is an altermagnet with a $d$-wave symmetry. The spin space group symmetry dictates a non-zero splitting with opposite chirality of the magnon bands along the diagonal directions of the Brillouin zone. This is confirmed by the LSW calculation shown in Fig. 1(c). Along the M-$\Gamma$-M direction, the acoustic (Goldstone) modes split into two branches with left- and right-handed chiralities, respectively. The splitting between the two chiral modes along $\Gamma$-M and $\Gamma$-M directions has opposite sign, and its magnitude reaches a maximum at the mid-point $\Sigma$ (and $\Sigma^{\prime}$). On the other hand, the two modes are degenerate along the $\Gamma$-X (and X-M) directions (Fig. 2(a)). These properties verify that the altermagnet has a $d$-wave nature. Note that the optical modes also split, but one band is completely flat along the M-$\Gamma$-M due to frustration and has zero spectral weight.
@@
Spin excitation spectra.—We implement tensor network calculation to go beyond the LSW results on the spin excitation spectra. The calculated spectrum at $J/J^{\prime}=1.1$ is shown in Fig. 2(b). It exhibits richer structures than the LSW results in Fig. 2(a). To understand these structures, we calculate the transverse and longitudinal components of the DSF, $\mathcal{S}^{xx+yy}(\mathbf{q},\omega)$ and $\mathcal{S}^{zz}(\mathbf{q},\omega)$, and show them in Fig. 2(c) and (d), respectively. The transverse DSF resembles the LSW spectrum with three dispersive magnon bands. The finite gap at M (and $\Gamma$) of the Goldstone modes is an effect of finite bond dimension $D$ of the iPEPS ansatz, and vanishes in the large-$D$ limit (Fig.S3 of SM SM ). One prominent feature is the splitting of Goldstone modes along the $\Gamma$-M direction. The opposite sign of $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$ for the two modes shown in Fig.S4 justifies their opposite chirality, as expected for an altermagnet.
@@
The spectrum of longitudinal modes, shown in Fig. 2(d), also consist of multiple bands, owing to the complex SS lattice structure (see discussion below). We can identify a mode (labeled by the dashed line) that develops a strong resonance at the M point. The energy of this mode decreases with reducing $J/J^{\prime}$ (see Fig.S2 of SM SM ). We then attribute it to the Higgs mode which reflects the amplitude fluctuations of the AFM order parameter. In a two-dimensional AFM state, the Higgs mode is usually hardly visible because the gapless Goldstone modes give rise to a divergent longitudinal susceptibility and the Higgs mode can decay into a pair of Goldstone modes Poldolsky_PRB84_174522_2011 . Here it shows up in the tensor network calculation for the following reasons: The gapped Goldstone modes at finite $D$ remove the divergence, and the single-mode approximation adopted suppresses the two-magnon continuum but preserves the two-magnon bound state, which is just the Higgs mode. Just above the Higgs mode, we find another longitudinal mode that is quasi-degenerate to the Higgs one at M point but develops a sizable separation at $\Sigma$ point. Similar behavior was reported in a related model Zhaojize_arxiv , and its origin needs further study. Besides the collective excitation modes discussed above, the spectrum also contains a continuum, as shown in the broad peak above about $2J^{\prime}$ in the density of states in Fig. S5 of SM SM . Interestingly, the peak moves to lower energy when reducing $J/J^{\prime}$ toward the plaquette-to-AFM transition, in a way similar to the behavior of the Higgs mode.
@@
Figure 3: (a): Spin DSF in the plaquette phase calculated by the iPEPS method with $D=3$ at $J/J^{\prime}=0.7$. The red dashed line labels the dispersion of the triplet excitations in the bond-operator theory. (b): $J/J^{\prime}$ dependence of extracted gaps at the M point of the Brillouin zone for several low-energy excitation modes in the plaquette phase, including, from bottom to top, a triplet, a singlet, and a 7-fold multiplet. The results are obtained by the iPEPS method with $D=4$. The dashed line shows the gap of the triplet excitations in the bond-operator theory.

### v2

# Spin excitations of the Shastry-Sutherland model – altermagnetism and deconfined quantum criticality

@@ Abstract
Frustrated quantum magnets can host a variety of exotic spin excitations, including fractionalized spin excitations coupled to emergent gauge fields at deconfined quantum critical points (DQCPs) and chiral magnons in altermagnets. Here, we investigate the spin excitation spectra of the highly frustrated $S=1/2$ antiferromagnetic (AFM) Shastry-Sutherland model, focusing on the evolution of low-energy collective modes from the Néel AFM phase to the plaquette valence bond solid (PVBS). We demonstrate that the AFM state exhibits altermagnetic behavior, characterized by a non-relativistic splitting between two chiral magnon bands. Furthermore, we identify two additional low-energy excitations: a Higgs mode in the longitudinal excitation channel and an $S=0$ excitation with vanishing spectral weight. As the system approaches the AFM-to-PVBS transition, both these modes soften along with the lowest-energy triplet and singlet modes in the PVBS state. The closing gap of the Higgs mode, combined with the nearly degenerate velocities of $S=1$ and $S=0$ excitations, provides spectral evidence that the AFM-to-PVBS transition is proximate to a DQCP with emergent $O(4)$ symmetry. Our results help clarify the spectral signature of a broad class of symmetry enhanced quantum phase transitions including deconfined quantum criticality.
I Introduction
@@
Symmetry provides an organizational principle to classify complex states of matter, and based on this, many exotic phenomena are understood TI_Review ; Senthil_AnnRevCondMattPhys_2015 ; Witten_NP_2018 . One prominent example is the recently proposed altermagnet, a collinear antiferromagnetic (AFM) state in which the two sublattices with oppositely polarized spins are connected by the rotational symmetry instead of inversion or translation Smejkal_PRX_2022A ; Smejkal_PRX_2022B . The distinct symmetry, which is organized by the spin space group theory Yang_arXiv:2105.12738 , causes a non-relativistic spin-dependent splitting of energy bands and may give rise to some exotic effects, such as anomalous Hall effect Smejkal_PRX_2022A ; Smejkal_NatRevMater7_2022 . Up to now, significant progress has been made in the study of itinerant systems Gonzalez_Betancourt_PRL_2023 ; Hariki_PRL_2024 ; Ferrari_arXiv:2408.00841 ; Wu_PRB_2007 ; Liu_NC_2021 ; Satoru_JPSJ_2019 . For insulating altermagnets, though a symmetry dictated splitting in magnon bands has been predicted Smejkal_PRL131_256703_2023 , existing studies are mostly limited to several model systems Ma_PRB110_064426_2024 ; Weissenhofer_PRB110_094427_2024 ; Yao_arXiv_2024 ; Zhaojize_arxiv .
@@
Symmetry can also dictate phase transitions. In the Landau-Ginzburg-Wilson paradigm, a system transitions from a disordered state to an ordered one by spontaneously breaking a symmetry. With quantum fluctuations the transition can be continuous and occur at zero temperature, giving rise to intriguing quantum critical phenomena sachdev_2011 ; Coldea_2010 ; Senthil2004 ; Cui2023DQCP ; Cui2019SCVO ; Lee2019 ; E8_2021 ; Xu2022 ; agrawal2020quantum ; fey2019quantum ; zhang2011exploring . For a transition between two ordered phases, the Landau paradigm predicts a first-order transition. However, this scenario has recently been challenged. It was proposed that the transition between a valence bond solid (VBS) and an antiferromagnetic (AFM) phase can be continuous, e.g., via a deconfined quantum critical point (DQCP) Senthil2004 . At the DQCP, the establishment of one order is accompanied by the destruction of the other order via proliferation of topological excitations, giving rise to emergent deconfined fractionalized excitations as well as enhanced continuous symmetry among the seemingly unrelated two types of order parameters. This scenario has motivated extensive studies on DQCP Sandvik2007 ; Meng_EasyplaneDQCP ; Shao_Science ; Senthil_Vojta_Sachdev:2004 ; Liu_Vojta:2022 ; Guo_PRL_2024 . However, whether a DQCP can be realized in quantum spin models and what would be its spectral signatures are still challenging questions.
@@
Tuning the frustration is a promising way to induce a VBS-to-AFM transition, and a prominent example is given by the Shastry-Sutherland (SS) model Shastry1981 . The model is defined on the SS lattice as illustrated in Fig. 1(a). It contains orthogonal spin dimers (with coupling $J^{\prime}$) connected by frustrated nearest neighbor (n.n.) bonds (with coupling $J$). With increasing $J/J^{\prime}$, the ground state experiences a series of quantum phase transitions from a dimer singlet (DS) phase to a plaquette valence bond solid (PVBS) at $J/J^{\prime}\approx 0.68$, then to a Néel AFM state for $J/J^{\prime}\sim 0.8$ (see Fig. 1(b)) Koga2000 ; Chung2001 ; Pixley2014 ; Corboz2013 ; Boos2019 ; Lee2019 ; yang2022quantum ; Sengupta_PRL:2013 ; Wang_PRL_2023 . Although the DS-to-PVBS transition is clearly strongly first-order, the nature of the PVBS-to-AFM transition remains elusive: Several tensor network calculations Corboz2013 ; Xi2023 showed it to be weakly first-order, while some other studies yang2022quantum ; Corboz_arXiv_2025 found an intervening gapless quantum spin liquid in between the PVBS and AFM phases. Interestingly, one DMRG work Lee2019 suggested the transition is via a DQCP with an emergent $O(4)$ symmetry, and quantum scaling behavior near this point has been revealed by a finite-size tensor network analysis WenyuanLiu_PRL_2024 . One prominent signature of a DQCP is the emergence of deconfined fractionalized spin excitations. Although spin excitations in both PVBS and AFM states have been extensively studied Mila_BOT ; Pinaki_PRB92_094440_2015 ; WangZ2018 ; Wang2023 , whether deconfined excitations emerge at the transition is still unclear.
@@
The phase diagram of the SS model well describes the evolution of low-temperature phases of the quasi-2D antiferromagnet $\mathrm{SrCu}_{2}(\mathrm{BO}_{3})_{2}$ under pressure tuning Kageyama1999 ; PhysRevLett.82.3701 ; Koga2000 ; Waki2007 ; haravifard2016crystallization ; Zayed2017 ; Bettler2020 ; Guo2020 ; Jimenez2021 , giving the hope to realize a DQCP in this material. Remarkably, a recent NMR study in this compound Cui2023DQCP reported an unusually large anomalous dimension, $\eta\sim 0.2$, at a field-induced PVBS-to-AFM transition under pressure, implying behavior in proximity to a DQCP. However, direct spectral evidence on the deconfined excitations at zero field is absent.
In this article, we study spin excitations of the SS model by using the state-of-the-art tensor network technique for infinite lattice systems, i.e., the infinite projected entangled pair state (iPEPS) method. This prevents the severe finite-size effect in usual numerical calculations and offers a new perspective on the nature of emergent DQCP phenomena. Rather than directly searching for the elusive emergent deconfined excitation continuum at the DQCP, we investigate the evolution of the collective excitation modes (namely, Higgs and Goldstone modes, etc.) associated with the broken emergent enhanced symmetry near the transition in the ordered phases. We demonstrate that their simultaneous, significant softening directly points to the existence of a DQCP. In the AFM state, we find that the altermagnetic nature of the Néel order results in a finite splitting with opposite chirality in the magnon bands along the $\Gamma$ to M direction of the Brillouin zone. Besides the transverse Goldstone excitations, we have identified two additional low-energy modes: a Higgs mode in the longitudinal excitation channel and an $S=0$ mode with vanishing spectral weight. As the system approaches the PVBS-to-AFM transition, these modes in the AFM phase, as well as the lowest triplet and singlet modes in the plaquette phase, soften and exhibit gap closing behavior at the transition point. We further find the velocities of the singlet, magnon, and Higgs modes become degenerate at the transition point. This, together with the gap closing behavior aforementioned, strongly supports the presence of a DQCP with an emergent O(4) symmetry at the PVBS-to-AFM transition.
@@
Figure 1: (a): Sketch of the SS lattice and the corresponding altermagnetism. The two sublattices of the Néel state are connected by a $C_{4}$ rotation of the lattice about the center of the empty plaquette (without the dimer bond). A time-reversal symmetry then recovers the Néel order. (b): Sketched phase diagram of the SS model with $J/J^{\prime}$. The $J/J^{\prime}$ values at the two transitions are taken from Ref. Xi2023 . (c): LSW dispersion in the AFM state of the SS model showing the chiral magnon bands with a non-zero splitting along the $\Gamma-M$ direction of the Brillouin zone. $\mathcal{L}$($\mathcal{R}$) refers to the left(right)-handed chiral magnon bands.
II Model and method
The Hamiltonian of the SS model reads as
@@ II Model and method
where $\mathbf{S}_{i}$ is an $S=1/2$ spin operator defined on site $i$, $J$ and $J^{\prime}$ are AFM inter- and intra-dimer couplings on the SS lattice, and we set $J^{\prime}=1$ in this work. We apply the iPEPS method verstraete2004renormalization ; orus2014practical ; corboz2014competing ; orus2009simulation ; PhysRevB.84.041108 to calculate the ground states of the model. The iPEPS wavefunction is obtained by variationally minimizing the ground state energy using gradient-based optimization, where the gradients are calculated through the automatic differentiation technique SRGAD ; liao2019differentiable . Then we construct the iPEPS ansatz for excited states through the single-mode approximation, and solve the generalized eigenvalue problem for the Hamiltonian in the tangent space spanned by these excited states  ponsioen2020excitations ; ponsioen2022automatic ; chi2022spin . Details of the methods are explained in the Appendix. With the information of the excited states, we calculate the spin dynamical structure factor (DSF) $\mathcal{S}^{\alpha\beta}(\mathbf{q},\omega)=\frac{1}{N}\sum_{i,j}\int dte^{i\mathbf{q}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}e^{i\omega t}\langle S_{i}^{\alpha}(t)S_{j}^{\beta}(0)\rangle$, where $\alpha,\beta$ refer to the spin components. The excitation spectra are calculated within the tensor network states up to bond dimension $D=5$. We also compare the results with linear spin wave (LSW) and bond-operator theories, which are introduced in detail in the Appendix.
III Altermagnetism and chiral magnons in the SS model
@@
The space group of the SS lattice is non-symmophic and the two sublattices of the Néel AFM state are connected by neither translational nor inversion, but a $C_{4}$ rotation about the center of the empty plaquette. The AFM order then recovers by further applying a time-reversal $\mathcal{T}$ symmetry to the spins (see Fig. 1(a)). This indicates that the Néel AFM state on the SS lattice is an altermagnet with a $d$-wave symmetry. The spin space group symmetry dictates a non-zero splitting with opposite chirality of the magnon bands along the diagonal directions of the Brillouin zone. This is confirmed by the LSW calculation shown in Fig. 1(c). Along the $\Gamma-M$ direction, the acoustic (Goldstone) modes split into two branches with left- and right-handed chiralities, respectively. The splitting between the two chiral modes along $\Gamma$-M and $\Gamma$-M directions has opposite sign, and its magnitude reaches a maximum at the mid-point $\Sigma$ (and $\Sigma^{\prime}$). On the other hand, the two modes are degenerate along the $\Gamma$-X (and X-M) directions (Fig. 2(a)). These properties verify that the altermagnet has a $d$-wave nature. Note that the optical modes also split, but one band is completely flat along the $\Gamma-M$ due to frustration and has zero spectral weight.
Figure 2: (a): Spin DSF in the AFM phase of the SS model calculated by the LSW theory. The inset shows high symmetric directions of the Brillouin zone. (b): Total spin DSF ($\mathcal{S}^{xx+yy+zz}$) in the AFM phase calculated by the iPEPS method with $D=3$ at $J/J^{\prime}=1.1$. (c): The transverse component ($\mathcal{S}^{xx+yy}$) of the results in (b). (d): The longitudinal component ($\mathcal{S}^{zz}$) of the results in (b). The red, black, and white lines with dots in (b) and (d) label the magnon, Higgs, and $S=0$ modes, respectively.
IV Spin excitation spectra
@@
We implement tensor network calculation to go beyond the LSW results on the spin excitation spectra. The calculated spectrum at $J/J^{\prime}=1.1$ is shown in Fig. 2(b). It exhibits richer structures than the LSW results in Fig. 2(a). To understand these structures, we calculate the transverse and longitudinal components of the DSF, $\mathcal{S}^{xx+yy}(\mathbf{q},\omega)$ and $\mathcal{S}^{zz}(\mathbf{q},\omega)$, and show them in Fig. 2(c) and (d), respectively. The transverse DSF resembles the LSW spectrum with three dispersive magnon bands. The finite gap at M (and $\Gamma$) of the Goldstone modes is an effect of finite bond dimension $D$ of the iPEPS ansatz, and vanishes in the large-$D$ limit (Fig.7 of the Appendix). One prominent feature is the splitting of Goldstone modes along the $\Gamma$-M direction. The opposite sign of $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$ for the two modes shown in Fig. 8 justifies their opposite chirality, as expected for an altermagnet.
@@
The spectrum of longitudinal modes, shown in Fig. 2(d), also consists of multiple bands, owing to the complex SS lattice structure (see discussion below). We can identify a mode (labeled by the black dashed line) that develops a strong resonance at the M point. The excitation gap of this mode decreases with reducing $J/J^{\prime}$ (see Fig.6 of the Appendix). We then attribute it to the Higgs mode which reflects the amplitude fluctuations of the AFM order parameter. In a two-dimensional AFM state, the Higgs mode is usually hardly visible because the gapless Goldstone modes give rise to a divergent longitudinal susceptibility and the Higgs mode can decay into a pair of Goldstone modes Poldolsky_PRB84_174522_2011 . Here it shows up in the tensor network calculation for the following reasons: The gapped Goldstone modes at finite $D$ remove the divergence, and the single-mode approximation adopted suppresses the two-magnon continuum but preserves the two-magnon bound state, which is just the Higgs mode. Just slightly above the Higgs mode, we find another excitation mode that is quasi-degenerate to the Higgs one at M point but with vanishing spectral weight. The gap of this mode follows a trend similar to that of the Higgs mode as $J/J^{\prime}$ decreases. We therefore identify this mode as a singlet ($S=0$) excitation originating from fluctuations of the plaquette VBS order parameter. Besides the collective excitation modes discussed above, the spectrum also contains a continuum, as shown in the broad peak above about $2J^{\prime}$ in the density of states in Fig. 9 of the Appendix. Interestingly, the peak moves to lower energy when reducing $J/J^{\prime}$ toward the PVBS-to-AFM transition, in a way similar to the behavior of the Higgs mode.
@@
Figure 3: (a): Spin DSF in the plaquette phase calculated by the iPEPS method with $D=3$ at $J/J^{\prime}=0.7$. The white line with dots shows the lowest singlet excitation. The magenta dashed line labels the dispersion of the triplet excitations in the bond-operator theory. (b): $J/J^{\prime}$ dependence of extracted gaps at the M point of the Brillouin zone for several low-energy excitation modes in the plaquette phase, including, from bottom to top, a triplet (in red), a singlet (in black), and a 7-fold multiplet (in blue). The results are obtained by the iPEPS method with $D=5$. The dashed line shows the gap of the triplet excitations in the bond-operator theory.